OpenStudy (anonymous):

The gasoline gauge on a van initially read 1/3 full. When 10 gallons were added to the tank, the gauge read 3/4 full. How many more gallons are needed to fill the tank? show work please:)

jimthompson5910 (jim_thompson5910):

Let x = total number of gallons "The gasoline gauge on a van initially read 1/3 full."----> the amount of gas in the tank at first is x/3 gallons "When 10 gallons were added to the tank, the gauge read 3/4 full" ---> x/3 + 10 = 3x/4 Now solve that equation for x x/3 + 10 = 3x/4 12*(x/3) + 12*10 = 12*(3x/4) 4x + 120 = 9x 120 = 9x - 4x 120 = 5x 120/5 = x 24 = x x = 24 So there are 24 gallons total. At first, there are x/3 = 24/3 = 8 gallons in the tank. You add 10 more to get 8+10 = 18 gallons (which is 3/4 of 24) So you still need 24 - 18 = 6 gallons